OpenStudy (anonymous):

what is 60/30 as a mixed number

OpenStudy (anonymous):

It's 2

OpenStudy (anonymous):

60÷30=2

OpenStudy (anonymous):

2 isn't a mixed number...

OpenStudy (anonymous):

well then you could do 1 30/30

OpenStudy (anonymous):

the question is 3 and 1/6 times 2/5 I used 30 as a common denominator and now I'm trying to figure out the answer

OpenStudy (anonymous):

oh that makes sense. thank you netertareh :)

OpenStudy (anonymous):

Problem as entered: 60 / 30 Divide the numerator, '60' by the denominator, '30', to get 2 with remainder 0. The remainder is the new numerator. The quotient, '2', becomes the whole number in the mixed number. 2 0 / 30 The resulting mixed number is: 2 0 / 30 Since the numerator in the mixed number is 0, the mixed number may be represented by just a whole number. 2 Disclaimer

OpenStudy (anonymous):

Bye people, I'm off to take my shower that I've been waiting for since 7 this morning!! *waves*

OpenStudy (texaschic101):

3 1/6 * 2/5 (turn to improper fraction) 19/6 * 2/5 = 38/30 = 1 8/30 reduces to 1 4/15 when you are multiplying fractions , you do not need a common denominator